OpenStudy (anonymous):

5. Explain, in complete sentences, how you would use the two methods (distribution and box method) to find the product of (x – 2) (x – 3). Please include your steps for each method and the resulting product. Which method do you prefer and why?

OpenStudy (anonymous):

This is what I have so far: I prefer the distribution method because, first of all, I find it easier, and secondly, that is how Mrs. Huerkamp explained it in the live lesson. To do this you simply utilize the FOIL technique by multiplying the first, outer, inner and last variables together. (x) (x) = x2 (x) (-3) = -3x (-2) (x) = -2x (-2) (-3) = 6 so your answer is: x2 - 3x - 2x + 6 = x2 - 5x + 6 The box method would be you make a 2x2 box and go up, down, and sideways to multiply the variables. It’s the same basic principal, just a little different: Now, I need help with the box method especially. idk how to write it out using the box method on the computer?? Can someone please help me!!

OpenStudy (anonymous):

"Hero" already answer this question, but he left out the box method part... This is what he wrote: It would be easier to say that for distribution method, recall that the formula for distributive property is a(b+c) = ab + bc. In this case: a = (x-2) b = x c = -3 Therefore: a(b+c) = ab + bc (x-2)(x-3) = (x-2)x + (x-2)-3 Thus: (x-2)(x-3) = (x-2)x + (x-2)-3 = x^2 - 2x - 3x + 6 = x^2-5x + 6

OpenStudy (anonymous):

Can anybody help me??

OpenStudy (anonymous):

wow, i am not familiar with the box method, let me look on line

OpenStudy (anonymous):

okay, thanks for looking at it a least... lol :)

OpenStudy (anonymous):

and neither am I so that's why its so hard for me, I know the distribution one pretty good, but i'm stumped when it comes to box method...

OpenStudy (anonymous):

at least* spelling error. hehehe :)

OpenStudy (anonymous):

Im not sure.. Sorry :/

OpenStudy (anonymous):

its okay, don't be sorry. It's not your fault, I just hate this freakin box method.... it sucks!

OpenStudy (anonymous):

I agree lol

OpenStudy (anonymous):

I don't even understand why the teacher wants me to write it in box method but whatever, I have to do it I one way or another.... lol :)

OpenStudy (accessdenied):

 seems like, you line up one binomial on top of 2x2 box and other binomial on side in descending degree, multiply together the corresponding row and column into each box, and add together i guess its just another way to do it, in case you preferred it that way